# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 7. Similarity
i.e., $(r,s)\in\varphi_{{\mathbf{B}}}$. Also recall that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\in D^{o}_{{\mathbf{B}}}$, and since $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$ we also get $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$. So $r,s\in D^{o}_{{\mathbf{B}}}$ and $(r,s)\in\varphi_{{\mathbf{B}}}$. But $D^{o}_{{\mathbf{B}}}$ is a transversal for $\varphi_{{\mathbf{B}}}$. These facts imply $r=s$, which implies $(a_{1},a_{1},b_{1},b_{1})\in T$. A similar argument gives $(a_{2},a_{2},b_{2},b_{2})\in T$. This proves (B3). Thus $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$.
(2) $\Rightarrow$ (4). Let ${\mathbf{C}}\in\mathcal{V}$, homomorphisms $f_{1}:{\mathbf{C}}\rightarrow{\mathbf{A}}$ and $f_{2}:{\mathbf{C}}\rightarrow{\mathbf{B}}$, and congruences $\delta_{1},\delta_{2},\delta^{+}_{1},\delta^{+}_{2},\psi,\tau$ satisfy the conditions in (2). Because $(\delta_{i},\delta^{+}_{i})\searrow(\psi,\tau)$, we get $\delta^{+}_{i}=\delta_{i}\circ\tau\circ\delta_{i}$ for $i=1,2$ and $(\delta_{1}:\delta^{+}_{1})=(\delta_{2}:\delta^{+}_{2})$ by Proposition 3.13.
Claim 7.8.1.
$f_{1}(\tau)=\mu$, $f_{2}(\tau)=\kappa$, and $f_{1}^{-1}(\alpha)=f_{2}^{-1}(\beta)$.
Proof of Claim 7.8.1.
We have $f_{1}(\tau)\subseteq\mu$ because $\tau\leq\delta^{+}_{1}$. Given $(a,b)\in\mu$, pick $c\in f_{1}^{-1}(a)$ and $d\in f_{1}^{-1}(b)$. Then $(c,d)\in\delta^{+}_{1}$, so there exist $c^{\prime},d^{\prime}\in C$ with
$c\stackrel{{\scriptstyle\delta_{1}}}{{\equiv}}c^{\prime}\stackrel{{\scriptstyle\tau}}{{\equiv}}d^{\prime}\stackrel{{\scriptstyle\delta_{1}}}{{\equiv}}d.$
Then $(a,b)=(f_{1}(c^{\prime}),f_{1}(d^{\prime}))\in f_{1}(\tau)$, proving $\mu\subseteq f_{1}(\tau)$. Thus $f_{1}(\tau)=\mu$, and a similar proof shows $f_{2}(\tau)=\kappa$. Finally, observe that
$f_{1}^{-1}(\alpha)=f_{1}^{-1}((0:\mu_{{\mathbf{A}}}))=(f_{1}^{-1}(0):f_{1}^{-1}(\mu_{{\mathbf{A}}}))=(\delta_{1}:\delta^{+}_{1}).$
A similar calculation shows $f_{2}^{-1}(\beta)=(\delta_{2}:\delta^{+}_{2})$. Since $(\delta_{1}:\delta^{+}_{1})=(\delta_{2}:\delta^{+}_{2}$), we get $f_{1}^{-1}(\alpha)=f_{2}^{-1}(\beta)$. ∎
Now define
$T=\{(f_{1}(c),f_{1}(d),f_{2}(c),f_{2}(d))\,:\,\text{$(c,d)\in\tau$}\}.$
We will show that $T$ is a proper bridge from ${\mathbf{A}}$ to ${\mathbf{B}}$. Clearly $T\leq{\mathbf{A}}\times{\mathbf{A}}\times{\mathbf{B}}\times{\mathbf{B}}$, and $\operatorname{pr}_{1,2}(T)=\mu$ and $\operatorname{pr}_{3,4}(\tau)=\kappa$ by Claim 7.8.1, proving (B1). $T$ clearly satisfies (B3), as $\tau$ is reflexive. To prove (B2), let $(a_{1},a_{2},b_{1},b_{2})\in T$ and pick $(c,d)\in\tau$ witnessing this. If $a_{1}=a_{2}$, then $(c,d)\in\ker(f_{1})=\delta_{1}$, so $(c,d)\in\tau\wedge\delta_{1}=\psi\leq\delta_{2}=\ker(f_{2})$, proving $b_{1}=b_{2}$. The opposite implication is proved similarly. ∎
@@
    is a proper bridge from ${\mathbf{A}}$ to $D({\mathbf{A}})$.
@@
(3) Since $(0:\mu)=\mu$, we have $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$ where $\Delta=\Delta_{\mu,\mu}$. Since $\mu$ is abelian, we get $\Delta=\left\{((a,b),(a^{\prime},b^{\prime})):\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}\in M(\mu,\mu)\right\}$ by Lemma 4.4(3). Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{A}}_{0}$ be the homomorphism with $\pi{\upharpoonright}_{A_{0}}=\mathrm{id}_{A_{0}}$ and $\ker(\pi)=\mu$. Then define $\lambda:{\mathbf{A}}\rightarrow D({\mathbf{A}})$ by $\lambda(x)=(x,\pi(x))/\Delta$. Suppose $(a,b)\in\ker(\lambda)$. Then $((a,\pi(a),(b,\pi(b)))\in\Delta$, so $M:=\begin{pmatrix}a&b\\ \pi(a)&\pi(b)\end{pmatrix}\in M(\mu,\mu)$. Also $(a,b)\in\mu$, so $\pi(a)=\pi(b)$. Thus the bottom entries of $M$ are equal. Since $\mu$ is abelian, it follows that the top entries of $M$ are also equal, i.e., $a=b$. Thus $\lambda$ is injective.
To prove that $\lambda$ is surjective, let $(a,b)\in\mu$ be given and define $e=\pi(a)=\pi(b)$ and $c=d(a,b,e)$ where $d(x,y,z)$ is a weak difference term for $\mathcal{V}$. Then $(a,b)/\Delta=(c,e)/\Delta=\lambda(c)$. Thus $\lambda$ is an isomorphism from ${\mathbf{A}}$ to $D({\mathbf{A}})$. ∎
Recognizing when two subdirectly irreducible algebras with abelian monoliths are similar is not always easy. Our final result of this section gives some “obvious” necessary conditions.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with abelian monoliths $\mu,\kappa$ respectively. Define $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$, let $\mathbb{F}_{{\mathbf{A}}}$ be the division ring associated to ${\mathbf{A}},\mu$ and some transversal for $\mu$, and let $\mathbb{F}_{{\mathbf{B}}}$ be the division ring associated to ${\mathbf{B}},\kappa$ and some transversal for $\kappa$. If ${\mathbf{A}}\sim{\mathbf{B}}$, then:
@@
    $\mathbb{F}_{{\mathbf{A}}}\cong\mathbb{F}_{{\mathbf{B}}}$.
@@
    If ${\mathbf{A}}$ and ${\mathbf{B}}$ are finite, then there exists an isomorphism $f:{\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$ such that for all $C\in A/\alpha$, $\max\{|e/\mu|\,:\,\text{$e\in C$}\}=\max\{|e^{\prime}/\kappa|\,:\,\text{$e^{\prime}\in f(C)$}\}$.
Proof.
Let $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$. Let $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ be the monoliths of $D({\mathbf{A}}),D({\mathbf{B}})$ respectively. Then $\lambda$ induces an isomorphism $\lambda^{\ast}:D({\mathbf{A}})/\varphi_{{\mathbf{A}}}\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ which moreover satisfies $|E|=|\lambda^{\ast}(E)|$ for every $\varphi_{{\mathbf{A}}}$-class $E$.
(1) ${\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and ${\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ by Corollary 4.8, so ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
(2) As in the proof of Theorem 7.8(1) $\Rightarrow$ (4) in the abelian monolith case, we can assume $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$. This implies that the division rings associated to $D({\mathbf{A}}),\varphi_{{\mathbf{A}}},D^{o}_{{\mathbf{A}}}$ and to $D({\mathbf{B}}),\varphi_{{\mathbf{B}}},D^{o}_{{\mathbf{B}}}$ are isomorphic. By Proposition 6.16, these division rings in turn are isomorphic to $\mathbb{F}_{{\mathbf{A}}}$ and $\mathbb{F}_{{\mathbf{B}}}$ respectively.
(3) Let $h_{{\mathbf{A}}}^{\ast}:{\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and $h_{{\mathbf{B}}}^{\ast}:{\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ be given by Corollary 4.8 Then for each $C\in A/\alpha$, $h_{{\mathbf{A}}}^{\ast}(C)=0_{C}/\varphi_{{\mathbf{A}}}$, which is the universe of $\operatorname{Grp}_{D({\mathbf{A}})}(\varphi_{{\mathbf{A}}},0_{C})$. Thus by Lemma 5.2(1) and Corollary 5.8(2),
$\max\{|e/\mu|\,:\,\text{$e\in C$}\}=|0_{C}/\varphi_{{\mathbf{A}}}|=|h^{\ast}_{{\mathbf{A}}}(C)|.$
Similarly, for any $\beta$-class $C^{\prime}$ we have
$\max\{|e^{\prime}/\kappa|\,:\,\text{$e^{\prime}\in C^{\prime}$}\}=|h^{\ast}_{{\mathbf{B}}}(C^{\prime})|.$
If $C^{\prime}\in B/\beta$ is chosen so that $h_{{\mathbf{B}}}^{\ast}(C^{\prime})=\lambda^{\ast}(h_{{\mathbf{A}}}^{\ast}(C))$, then $|h^{\ast}_{{\mathbf{B}}}(C^{\prime})|=|h^{\ast}_{{\mathbf{A}}}(C)|$ by a comment in the preamble to this proof. Thus we can set $f=(h^{\ast}_{{\mathbf{B}}})^{-1}\circ\lambda^{\ast}\circ h^{\ast}_{{\mathbf{A}}}$. ∎
8. Examples
In this section we illustrate some of our earlier results by presenting a recipe for constructing finite subdirectly irreducible algebras with abelian monoliths in a varieties having weak difference terms. The algebras ${\mathbf{A}}$ can be constructed so that if $\mu$ is its monolith and $\alpha=(0:\mu)$, then

### v2

@@ 7. Similarity
i.e., $(r,s)\in\varphi_{{\mathbf{B}}}$. Also recall that $(a_{1},a_{1})/\Delta_{{\mathbf{A}}}\in D^{o}_{{\mathbf{A}}}$ and $s=(b_{1},b_{1})/\Delta_{{\mathbf{B}}}\in D^{o}_{{\mathbf{B}}}$, and since $\lambda(D^{o}_{{\mathbf{A}}})=D^{o}_{{\mathbf{B}}}$ we also get $r=\lambda((a_{1},a_{1})/\Delta_{{\mathbf{A}}})\in D^{o}_{{\mathbf{B}}}$. So $r,s\in D^{o}_{{\mathbf{B}}}$ and $(r,s)\in\varphi_{{\mathbf{B}}}$. But $D^{o}_{{\mathbf{B}}}$ is a transversal for $\varphi_{{\mathbf{B}}}$. These facts imply $r=s$, which implies $(a_{1},a_{1},b_{1},b_{1})\in T$. A similar argument gives $(a_{2},a_{2},b_{2},b_{2})\in T$. This proves (B3). Thus $T$ is a similarity bridge from $\mathbf{A}$} to $\mathbf{B}$}. ∎
@@
    is a similarity bridge from $\mathbf{A}$} to $D({\mathbf{A}})$.
@@
(3) Since $(0:\mu)=\mu$, we have $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$ where $\Delta=\Delta_{\mu,\mu}$. Since $\mu$ is abelian, we get $\Delta=\left\{((a,b),(a^{\prime},b^{\prime})):{\scriptsize{\begin{pmatrix}a&a^{\prime}\\ b&b^{\prime}\end{pmatrix}}}\in M(\mu,\mu)\right\}$ by Lemma 4.4(4). Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{A}}_{0}$ be the homomorphism with $\pi{\upharpoonright}_{A_{0}}=\mathrm{id}_{A_{0}}$ and $\ker(\pi)=\mu$. Then define $\lambda:{\mathbf{A}}\rightarrow D({\mathbf{A}})$ by $\lambda(x)=(x,\pi(x))/\Delta$. Suppose $(a,b)\in\ker(\lambda)$. Then $((a,\pi(a),(b,\pi(b)))\in\Delta$, so $M:={\scriptsize{\begin{pmatrix}a&b\\ \pi(a)&\pi(b)\end{pmatrix}\in M(\mu,\mu)}}$. Also $(a,b)\in\mu$, so $\pi(a)=\pi(b)$. Thus the bottom entries of $M$ are equal. Since $\mu$ is abelian, it follows that the top entries of $M$ are also equal, i.e., $a=b$. Thus $\lambda$ is injective.
To prove that $\lambda$ is surjective, let $(a,b)\in\mu$ be given and define $e=\pi(a)=\pi(b)$ and $c=d(a,b,e)$ where $d(x,y,z)$ is a weak difference term for $\mathcal{V}$. Then $(a,b)/\Delta=(c,e)/\Delta=\lambda(c)$ as required. ∎
We end this section with two easy consequences of similarity in the abelian monolith case.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with abelian monoliths $\mu,\kappa$ respectively. Define $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$, and let $\mathbb{F}_{\mu}$ and $\mathbb{F}_{\kappa}$ be the division rings associated to $({\mathbf{A}},\mu)$ and $({\mathbf{B}},\kappa)$ respectively. If ${\mathbf{A}}\sim{\mathbf{B}}$, then:
@@
    $\mathbb{F}_{\mu}\cong\mathbb{F}_{\kappa}$.
Proof.
(1) Let $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$. Let $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ be the monoliths of $D({\mathbf{A}}),D({\mathbf{B}})$ respectively. Then $\lambda$ induces an isomorphism $\lambda^{\ast}:D({\mathbf{A}})/\varphi_{{\mathbf{A}}}\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$. We have ${\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and ${\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ by Corollary 4.8, so ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
(2) follows from Proposition 6.9. ∎
8. Conclusion and questions
In this paper we have extended Freese’s analysis of abelian minimal congruences, and his similarity relation between subdirectly irreducible algebras, from the congruence modular setting to the setting of varieties with a weak difference term. We also established a connection between similarity and certain constructs which we have called “similarity bridges.” In the companion paper [31], we will compare similarity bridges to Zhuk’s original bridges (which are superficially weaker) and explain some related elements of Zhuk’s work in terms of centrality and similarity. In a third paper [30], we will unite the analyses of “rectangular critical relations” by Zhuk [35, Lemma 8.21] (in the idempotent locally finite Taylor case) and by Kearnes and Szendrei [19] (in the congruence modular case), showing in particular the role of similarity, associated finite fields, and linear equations.
Problem 1.
Finite subdirectly irreducible algebras with abelian monoliths in idempotent locally finite Taylor varieties are of particular interest. We have seen one result (Lemma 5.9) showing how the assumption of idempotency can restrict the possible sizes of monolith classes. In what other ways does idempotency affect the sizes of monolith classes?
Problem 2.
Suppose $\mathbf{A}$} is a finite idempotent subdirectly irreducible algebra in a Taylor variety with abelian monolith $\mu$, $\alpha=(0:\mu)$, and $E$ is an $\alpha$-class. If $q$ is the maximal size of a $\mu$-class in $E$, does it follow that every $\mu$-class in $E$ has size $q$ or 1?
When $\mathbf{A}$} is a finite algebra in a variety with a weak difference term and $\theta$ is an abelian minimal congruence, tame congruence theory tells us that $\operatorname{typ}_{{\mathbf{A}}}(0,\theta)=\mathbf{2}$ and therefore the minimal algebras ${\mathbf{A}}|_{N}$ associated to $\langle 0,\theta\rangle$-traces are weakly isomorphic to a one-dimensional vector space over a finite field $\mathbb{F}$ determined by $\theta$.
Problem 3.
Is the finite field $\mathbb{F}$ mentioned above isomorphic to $\mathbb{F}_{\theta}$? If not, how are $\mathbb{F}$ and $\mathbb{F}_{\theta}$ related?
Appendix 1: One division ring
Let $\mathbf{A}$} be an algebra in a variety with a weak difference term, let $\theta$ be an abelian minimal congruence, and let $\mathbb{F}_{\theta}$ be the division ring constructed in Section 6. In this appendix we will show that $\mathbb{F}_{\theta}$ is isomorphic to another division ring $\mathbb{D}$ associated to $\theta$, which Freese [5, Section 4, page 150] described in the congruence modular setting and whose construction easily adapts to algebras having a weak difference term. We also show that when $\mathbf{A}$} is finite, $\mathbb{F}_{\theta}$ is isomorphic to the finite field $\mathbb{F}$ arising in the proof sketch of Theorem 3.9.
First we describe our adaptation of Freese’s construction. Assume only that $\mathbf{A}$} is an algebra having a weak difference term $d$, and $\theta$ is an abelian minimal congruence. Let the $\theta$-classes be indexed by a set $I$ via the notation $A/\theta=\{C_{i}\,:\,\text{$i\in I$}\}$. Let $T$ be a transversal for $\theta$ and for each $i\in I$ let $e_{i}$ be the unique element of $T\cap C_{i}$. For $i,j\in I$ define
$R_{ij}=\{f{\upharpoonright}_{C_{j}}\,:\,\text{$f\in\operatorname{Pol}_{1}({\mathbf{A}})$ and $f(e_{j})=e_{i}$}\}.$
(Freese denoted this set by $\operatorname{Hom}(\theta,e_{j},e_{i})$.) Also define $O_{ij}\in R_{ij}$ to be the constant map $C_{j}\rightarrow\{e_{i}\}$. Let $\mathcal{R}$ be the set of $I\times I$-matrices $(r_{ij})_{i,j\in I}$ where $r_{ij}\in R_{ij}$ for all $i,j\in I$, and for each $j\in I$, $r_{ij}=O_{ij}$ for all but finitely many $i\in I$. Let $V$ denote the direct sum of the abelian groups $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$, $i\in I$, realized as the subgroup of $\prod_{i\in I}\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ consisting of those $(a_{i})_{i\in I}\in\prod_{i\in I}C_{i}$ for which $a_{j}=e_{j}$ for all but finitely many $j\in I$. Observe that each $f\in R_{ij}$ is a group homomorphism from $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{j})$ to $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$, by Proposition 3.8; that $R_{ij}$ is closed under the pointwise addition operation in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$; and that $f\in R_{ij}$, $g\in R_{jk}$ imply $f\circ g\in R_{ik}$.
Thus $\mathcal{R}$ acts in the usual “matrix” way on $V$: given $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$, define $\mathsf{L}_{M}:V\rightarrow V$ so that $\mathsf{L}_{M}((a_{i})_{i\in I})=(b_{i})_{i\in I}$ where
$b_{i}=\sum_{j\in I,\,a_{j}\neq e_{j}}r_{ij}(a_{j})\quad\mbox{with the sum computed in $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$}.$
One can show that each $\mathsf{L}_{M}\in\operatorname{End}(V)$, the map $M\mapsto\mathsf{L}_{M}$ is injective, and its range $\mathsf{L}_{\mathcal{R}}:=\{\mathsf{L}_{M}\,:\,\text{$M\in\mathcal{R}$}\}$ is a unital subring of $\mathbf{End}(V)$. Via the identification of $\mathcal{R}$ with $\mathsf{L}_{\mathcal{R}}$, $\mathcal{R}$ inherits the operations of a unital ring, and we obtain the faithful left $\mathcal{R}$-module $_{\mathcal{R}}V$. The minimality of $\theta$ implies that $_{\mathcal{R}}V$ is a simple module. Freese [5, p. 150] explained this in the congruence modular case; here is a proof that does not require congruence modularity. Let $\overline{0}=(e_{i})_{i\in I}\in V$. The claim of simplicity reduces to proving the following: if $\mathbf{a}=(a_{i})_{i\in I}\in V$ with $\mathbf{a}\neq\overline{0}$, and $\mathbf{b}=(b_{i})_{i\in I}\in V$ is such that there exists $k\in I$ with $b_{i}=e_{i}$ for all $i\in I\setminus\{k\}$, then $\mathbf{b}\in\mathcal{R}\mathbf{a}$. To prove it, pick $\ell\in I$ with $a_{\ell}\neq e_{\ell}$. By minimality of $\theta$ we have $(b_{k},e_{k})\in\operatorname{Cg}(a_{\ell},e_{\ell})$. Thus by Corollary 3.2, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(a_{\ell})=b_{k}$ and $f(e_{\ell})=e_{k}$. Let $M=(r_{ij})_{i,j\in I}\in\mathcal{R}$ be the matrix given by $r_{k\ell}=f{\upharpoonright}_{C_{\ell}}$ and $r_{ij}=O_{ij}$ for $(i,j)\neq(k,\ell)$. Then $M\mathbf{a}=\mathsf{L}_{M}(\mathbf{a})=\mathbf{b}$ as required. It follows by Schur’s Lemma for modules that the endomorphism ring $\mathbf{End}({}_{\mathcal{R}}V)$ of the module $_{\mathcal{R}}V$ is a division ring $\mathbb{D}$. This is the division ring Freese associated to $\theta$ and the transversal $T$ (in the congruence modular case).
Proposition A1.1.
Suppose $\mathbf{A}$} is an algebra in a variety with a weak difference term and $\theta$ is an abelian minimal congruence of $\mathbf{A}$}. Let $\mathbb{D}$ be the division ring associated to $\theta$ and some transversal $T$ in the previous two paragraphs. Then $\mathbb{D}\cong\mathbb{F}_{\theta}$.
Proof sketch.
Let $\alpha=(0:\theta)$ and $\Delta=\Delta_{\theta,\alpha}$. Recall that the universe of $\mathbb{F}_{\theta}$ is the set of all endomorphisms $\lambda$ of the difference algebra $D({\mathbf{A}},\theta)={\mathbf{A}}(\theta)/\Delta$ which preserve $\varphi:=\overline{\alpha}/\Delta$ and fix the elements of $D^{o}=\{0_{E}\,:\,\text{$E\in A/\alpha$}\}$ pointwise. Let $V$ be the abelian group defined in the paragraphs preceding Proposition A1.1. Define the ring ${\mathbf{P}}=\prod_{i\in I}\mathbf{End}(\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i}))$ and let $\Psi:{\mathbf{P}}\rightarrow\mathbf{End}(V)$ be the obvious (coordinatewise) ring embedding. Given $\lambda\in\mathbb{F}_{\theta}$, define $\Phi(\lambda)=(\lambda_{i})_{i\in I}$ where for each $i\in I$, $\lambda_{i}:C_{i}\rightarrow C_{i}$ is the map given as follows: given $a\in C_{i}$, $\lambda_{i}(a)$ is the unique $b\in C_{i}$ such that $\lambda((a,e_{i})/\Delta)=(b,e_{i})/\Delta$ (see Corollary 6.7). It is tedious but straightforward to check that $\Phi$ is an injective ring homomorphism from $\mathbb{F}_{\theta}$ into $\mathbf{P}$}.
It remains only to show that the range of $\Psi\circ\Phi$ is precisely $\operatorname{End}({}_{\mathcal{R}}V)$, for then $\Psi\circ\Phi$ will be an isomorphism from $\mathbb{F}_{\theta}$ to $\mathbb{D}$. We can establish this by proving the following:
@@ 8. Conclusion and questions
    $\operatorname{End}({}_{\mathcal{R}}V)\subseteq\Psi(P)$.